40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 65 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  13720   Fri Mar 30 03:23:50 2018 KojiUpdateGeneralaLIGO EOM work

I have been working on the aux beat setup on the PSL table between 9PM-3AM.

This work involved:

- Turning off the main marconi
- Turning off the freq generation unit (incl IMC modulation)
- Closing the PSL shutter

After the work, these were reverted and the IMC and both arms have been locked.

  13719   Thu Mar 29 17:57:36 2018 arijitUpdateIOOMCREFL_PD Optical response measurement

Kevin, Gautam and Arijit

Today we performed the in-loop noise measurements of the MCREFL-PD using the SR785 to ascertain the effect of the Noise Eater on the laser. We took the measurements at the demodulated output channel from the MCREFL-PD. We performed a series of 6 measurements with the Noise Eater ''ON'' and ''OFF''. The first data set is an outlier probably, due to some transient effects. The remaining data sets were recorded in succession with a time interval of 5 minutes each between the Noise Eater in the ''ON'' and ''OFF'' state. We used the calibration factor of 13kHz/Vrms from elog 13696 to convert the V_rms to Hz-scale.

The conclusion is that the NOISE EATER does not have any noticeable effect on the noise measurements.

ALS beat spectrum and also the arm control signal look as they did before. coherence between arm control signals (in POX/POY lock) is high between 10-100Hz, so looks like there is still excess frequency noise in the MC transmitted light. Looking at POX as an OOL sensor with the arm under ALS control shows ~10x the noise at 100 Hz compared to the "nominal" level, consistent with what Koji and I observed ~3weeks ago.

We tried swapping out Marconis. Problem persists. So Marconi is not to blame. I wanted to rule this out as in Jan, Steve and I had installed a 10MHz reference to the rear of the Marconi.

Attachment 1: NOISE_EATER_On_OFF.pdf
NOISE_EATER_On_OFF.pdf
  13718   Thu Mar 29 17:14:42 2018 KiraUpdatePEMPID test

[Kira, Gautam]

We closed the loop today and implemented the PID script. I have attached the StripTool graph for an integral value of 0.5 and proportional value of 20. We had some issues getting it to work properly and it would oscillate between some low values of the control voltage. The set point here was -3.20, which corresponds to about a 20 degree increase in temperature. The next step would be to find which values of Kp, Ki, and Kd would work in this case and low pass filter the signal from the temperature sensor, and also create an MEDM screen for easier PID control.

Attachment 1: PID_test.png
PID_test.png
  13717   Thu Mar 29 12:03:37 2018 Jon RichardsonSummaryGeneralProof-of-Concept SRC Gouy Phase Measurement

I've been developing an idea for making a direct measurement of the SRC Gouy phase at RF. It's a very different approach from what has been tried before. Prior to attempting this at the sites, I'm interested in making a proof-of-concept measurement demonstrating the technique on the 40m. The finesse of the 40m SRC will be slightly higher than at the sites due to its lower-transmission SRM. Thus if this technique does not work at the 40m, it almost certainly will not work at the sites.

The idea is, with the IFO locked in a signal-recycled Michelson configuration (PRM and both ETMs misaligned), to inject an auxiliary laser from the AS port and measure its reflection from the SRC using one of the pre-OMC pickoff RFPDs. At the sites, this auxiliary beam is provided by the newly-installed squeezer laser. Prior to injection, an AM sideband is imprinted on the auxiliary beam using an AOM and polarizer. The sinusoidal AOM drive signal is provided by a network analyzer, which sweeps in frequency across the MHz band and demodulates the PD signal in-phase to make an RF transfer function measurement. At the FSR, there will be a AM transmission resonance (reflection minimum). If HOMs are also present (created by either partially occluding or misaligning the injection beam), they too will generate transmission resonances, but at a frequency shift proportional to the Gouy phase. For the theoretical 19 deg one-way Gouy phase at the sites, this mode spacing is approximately 300 kHz. If the transmission resonances of two or more modes can be simultaneously measured, their frequency separation will provide a direct measurement of the SRC Gouy phase.

The above figure illustrates this measurement configuration. An attached PDF gives more detail and the expected response based on Finesse modeling of this IFO configuration.

Attachment 1: src_gouy_phase_v3.pdf
src_gouy_phase_v3.pdf src_gouy_phase_v3.pdf src_gouy_phase_v3.pdf src_gouy_phase_v3.pdf
  13716   Wed Mar 28 21:47:37 2018 arijitUpdateIOOMCREFL_PD Optical response measurement

Kevin, Gautam and Arijit

We did a optical measurement of the MCREFL_PD transimpedance using the Jenny Laser set-up. We used 0.56mW @1064nm on the NewFocus 1611 Photodiode as reference and 0.475mW @1064nm on the MCREFL_PD. Transfer function was measured using the AG4395 network analyzer. We also fit the data using the refined LISO model. From the optical measurement, we can see that we do not have a prominent peak at about 300MHz like the one we had from the electrical transimpedence measurement. We also put in the electrical transimpedence measurement as reference. RMS contribution of 300MHz peak to follow.

 

As per Rana`s advice I have updated the entry with information on the LISO fit quality and parameters used. I have put all the relevant files concerning the above measurement as well as the LISO fit and output files as a zip file "LISO_fit" . I also added a note describing what each file represents. I have also updated the plot with fit parameters and errors as in elog 10406.

Attachment 1: LISO_fit_with_info.pdf
LISO_fit_with_info.pdf
Attachment 2: LISO_fit.zip
  13715   Wed Mar 28 21:31:39 2018 gautamUpdateIOOMC REFL PD removed

I re-installed the MC REFL photodiode. Centered beam on the PD by adjusting steering mirror to maximize the DC signal level (on o'scope) at the DC monitoring port. Curiously, the DC level on the scope (high-Z) was ~2.66V DC, whereas the MEDM screen reports ~twice that value, at ~5.44 "V". We may want to fix this "calibration" (or better yet, use physical units like mW). Noise-eater On/Off comparison of MC error signals to follow.

  13714   Wed Mar 28 17:28:58 2018 SteveUpdatePSLnoise eater on or off

Till RIN measurement noise eater is off on 2W laser. The inno 1W  has no noise eater.

2010 power v current table is below.

Quote:

Koji and Kevin measured the output power vs injection current for the Innolight 2W laser.

The threshold current is 0.75 A.

 

The following data was taken with the laser crystal temperature at 25.04ºC (dial setting: 0.12).

Injection Current (A) Dial Setting Output Power (mW)
0.000 0.0 1.2
0.744 3.66 1.1
0.753 3.72 4.6
0.851 4.22 102
0.954 4.74 219
1.051 5.22 355
1.151 5.71 512
1.249 6.18 692
1.350 6.64 901
1.451 7.08 1118
1.556 7.52 1352
1.654 7.92 1546
1.761 8.32 1720
1.853 8.67 1855
1.959 9.05 1989
2.098 9.50 2146

 

 

Attachment 1: inno2W.png
inno2W.png
Attachment 2: inno1W.png
inno1W.png
  13713   Wed Mar 28 16:44:27 2018 SteveUpdateGeneralAP table today

MCRefl is absent, it is under investigation. I removed a bunch of hardware and note all spare optics along the edges.

 

Attachment 1: AP_Table_20180328.png
AP_Table_20180328.png
  13712   Tue Mar 27 23:37:35 2018 gautamUpdateIOOMC REFL PD removed

I've removed the MC REFL PD unit from the AS table for investigation. So MC won't lock.

PSL shutter was closed and location of PD was marked with sharpie (placing guides to indicate position wasn't convenient). I also kapton taped the PD to minimize dust settling on the PD while I have it out in the electronics area. Johannes has the camera, and my cellphone image probably isn't really high-res enough for diagnostics but I'm posting it here anyways for what it's worth. More importantly - the board is a D980454 revision B judging by the board, but there is no schematic for this revision on the DCC. The closest I can find is a D980454 Rev D. But I can already see several differences in the component layout (though not all of them may be important). Making a marked up schematic is going to be a pain indecision. I'm also not sure what the specific make of the PD installed is.

The lid of the RF cage wasn't on.

More to follow tomorrow, PD is on the electronics workmench for now...


gautam 28 March 2018: Schematic has been found from secret Dale stash (which exists in addition to the secret Jay stash). It has also been added to the 40m electronics tree.

Attachment 1: IMG_6955.JPG
IMG_6955.JPG
  13711   Tue Mar 27 19:32:03 2018 arijitUpdateIOOPSL noise eater was off

Kevin, Gautam and Arijit

We made a measurement of the MC_REFL photodiode transfer function using the network analyzer. We did it for two different power input (0dB and -10dB) to the test measurement point of the MC_REFL photodiode. This was important to ensure the measurements of the transfer function of the MC_REFL photodiode was in the linear regime. The measurements are shown in attachment 1. We corrected for phase noise for the length of cable (50cm) used for the measurement. With reference to ELOG 10406, in comparison to the transimpedance measurement performed by Riju and Koji, there is a much stronger peak around 290MHz as observed by our measurement.


We also did a noise measurement for the MC_REFL photodiode. We did it for three scenarios: 1. Without any light falling on the photodiode 2. With light falling on the photodiode, the MC misaligned and the NPRO noise eater was OFF 3. With light falling on the photodiode, the MC misaligned and the NPRO noise eater was ON. We observed that the noise eater does reduce the noise being observed from 80kHz to 20MHz. This is shown in attachment 2.


We did the noise modelling of the MC_REFL photodiode using LISO and tried matching the expected noise from the model with the noise measurements we made earlier. The modeled noise is in good agreement with the measured noise with 10Ohms in the output resistance. The schematic for the MC_REFL photodiode however reveals a 50Ohm resistance being used. The measured noise shows excess noise ~ 290MHz. This is not predicted from the simplied LISO model of the photodiode we took.


Discussion with Koji and Gautam revealed that we do not have the exact circuit diagram for the MC_REFL photodiode. Hence the simplified model that was assumed earlier is not able to predict the excess noise at high frequencies. One thing to note however, is that the excess noise is measured with the same amplitude even with no light falling on the MC_REFL photodiode. This means that there is a positive feedback and oscillation in the op-amp (MAX4107) at high frequencies. One way to refine the LISO model would be to physically examine the photodiode circuit.

We also recorded the POX and POY RF monitor photodiode outputs when the interferometer arms are independently stabilized to the laser. Given the noise outputs from the RF photodiodes were similar, we have only plotted the POY RF monitor output for the sake of clarity and convenience.

Quote:

While Kevin and Arijit were doing their MC_REFL PD noise measurements (which they will elog about separately shortly), I noticed a feature around 600kHz that reminded me of the NPRO noise eater feature. This is supposed to suppress the relaxation oscillation induced peak in the RIN of the PSL. Surprisingly, the noise eater switch on the NPRO front panel was set to "OFF". Is this the normal operating state? I thought we want the noise eater to be "ON"? Have to measure the RIN on the PSL table itself with one of the many available pick off PDs. In any case, as Attachment #1 showed, turning the noise eater back on did not improve the excess IMC frequency noise.

 

Attachment 1: MCREFL_TF.pdf
MCREFL_TF.pdf
Attachment 2: MCREFL_SPECTRUM.pdf
MCREFL_SPECTRUM.pdf
  13710   Tue Mar 27 11:11:16 2018 KiraUpdatePEMChannel setup

[Kira, Gautam]

We setup the channels for PID control of the seismometer can. First, we ssh into c1auxex and went to /cvs/cds/caltech/target/c1auxex2 and found ETMXaux.db. We then added in new soft channels that we named C1:PEM-SEIS_EX_TEMP_SLOWKP, C1:PEM-SEIS_EX_TEMP_SLOWKI, C1:PEM-SEIS_EX_TEMP_SLOWKD that will control the proportional, integral and differential gain respectively. These channels are used in the script FSSSlow.py for PID control. We then had to restart the system, but first we turned off the LSC mode and then shut down the watchdog on the X end. After doing the restart, we disabled the OPLEV as well before restarting the watchdog. Then, we enabled the LSC mode again. This is done to not damage any of the optics during the restart. The restart is done by using sudo systemctl restart modbusIOC.service and restarted with sudo systemctl status modbusIOC.service. Then, we made sure that the channels existed and could be read and writtten to, so we tried z read [channel name] and it read 0.0. We then did z write [channel name] 5, and it wrote 5 to that channel. Now that the channels work, we can implement the PID script and check to make sure that it works as well.

  13709   Tue Mar 27 08:58:21 2018 SteveUpdateVACVM1 opened.......scan fine

 

Quote:

CC1 old MKS cold cathode gauge randomly turns on- off. This makes software interlock close VM1 to protect RGA  So the closed off RGA region pressure goes up and the result is distorted RGA scan.

CC1 MKS gauge is disconnected and VM1 opened. This reminds me that we should connect our interlocks to CC1 Hornet Pressure gauge.

Quote:

Pumpdown 80 at 511 days and pd80b at 218 days

Valve configuration:  special vacuum normal, annuloses are not pumped at 3 Torr, IFO pressure 7.4e-6 Torr at vac envelope temp 22 +- 1C degrres

 

 

 

Attachment 1: rga2018march27.png
rga2018march27.png
  13708   Tue Mar 27 01:39:44 2018 gautamUpdateIOOPSL noise eater was off

While Kevin and Arijit were doing their MC_REFL PD noise measurements (which they will elog about separately shortly), I noticed a feature around 600kHz that reminded me of the NPRO noise eater feature. This is supposed to suppress the relaxation oscillation induced peak in the RIN of the PSL. Surprisingly, the noise eater switch on the NPRO front panel was set to "OFF". Is this the normal operating state? I thought we want the noise eater to be "ON"? Have to measure the RIN on the PSL table itself with one of the many available pick off PDs. In any case, as Attachment #1 showed, turning the noise eater back on did not improve the excess IMC frequency noise.

Attachment 1: IR_ALS_20180326.pdf
IR_ALS_20180326.pdf
  13707   Mon Mar 26 23:49:27 2018 gautamUpdateGeneralNew ADC Adaptor Board installed in C1LSC expansion chassis

Todd informed me that the ADC Timing adaptor boards we had ordered arrived today. I had to solder on the components and connectors as per the schematic, though the main labor was in soldering the high density connectors. I then proceeded to shut down all models on c1lsc (and then the FE itself). Then classic problem of all vertex machines crashing when unloading models on c1lsc happened (actually Koji noticed that this was happening even on c1ioo). Anyways this was nothing new so I decided to push ahead. 

I had to get a cable from Downs that connects the actual GS ADC card to this adaptor board. I powered off the expansion chassis, installed the adaptor board, connected it to the ADC card and restarted the expansion chassis and also the FE. I also reconnected the SCSI cable from the AA board to the adaptor card. It was a bit of a struggle to get all the models back up and running again, but everything eventually came back(after a few rounds of hard rebooting). I then edited the c1x04 and c1lsc simulink models to reflect the new path for the X arm ALS error signals. Seems to work alright.

At some point in the afternoon, I noticed a burning smell concentrated near the PSL table. Koji traced the smell down to the c1lsc expansion chassis. We immediately powered the chassis off. But Steve later informed me that he had already noticed an odd burning smell in the morning, before I had done any work at the LSC rack. Looking at the newly installed adaptor card, there wasn't any visual evidence of burning. So I decided to push ahead and try and reboot all models. Everything came back up normally eventually, see Attachment #1. Particle count in the lab seems a little higher than usual (actually, according to my midnight measurement, they are ~factor of 10 lower than Steve's 8am measurements), but Steve didn't seem to think we should read too much into this. Let's monitor the situation over the coming days, Steve should comment on the large variance seen in the particle counter output which seems to span 2 orders of magnitude depending on the time of the day the measurement is made... Also note that there is a BIO card in the C1LSC expansion chassis that is powered by a lab power supply unit. It draws 0 current, even though the label on it says otherwise. I a not sure if the observed current draw is in line with expectations.


The spare (unstuffed) adaptor cards we ordered, along with the necessary hardware to stuff them, are in the Digital FE hardware cabinet along the east arm.

Steve:  particle count in the 40m is following outside count, wind direction, weather condition .....etc. The lab particle count is NOT logged ! This is bad practice.

Attachment 1: CDS_20180326.png
CDS_20180326.png
  13706   Mon Mar 26 21:40:26 2018 gautamSummaryIOOMC2 classical radiation pressure noise

[rana, gautam]

we measured the RIN of the MC2 transmission using the PDA255 I had put on the MC2 trans table sometime ago for ringdowns. Attached are (i) spectra for the RIN, (ii) spectra for the classical rad. pressure noise assuming 500W circulating power and (iii) a tarball of data and code used to generate these plots.

We took a full span measurement (to make sure there aren't any funky high-freq features) and a measurement from DC-800 Hz (where we are looking for excess noise). The DC level of light on the photodiode was 2.76V (measured using o'scope)

I'll add this to the noise budget later. But the measured RIN seems consistent with a 2013 measurement at 100Hz (though the 2013 measurement is using DTT and so doesn't have high frequency information).

Attachment 1: IMC_RIN.pdf
IMC_RIN.pdf
Attachment 2: IMC_RadPress.pdf
IMC_RadPress.pdf
Attachment 3: MC2_radPress.tgz
  13705   Mon Mar 26 21:25:55 2018 ranaSummaryIOOMC2 Trans table has issues

Gautam, Rana

While at the MC2 table, we noticed that it has some optical problems:

  1. There is an ND filter mounted to the beam reducing lens. ND filters are illegal, Steve. It causes too much scattering noise. We should instead have a beamsplitter and a dump.
  2. One of those bad U100-AC28 mounts is in use. This is one of those ones with plastic clips that Osamu liked, but the plastic gets in the way of the beam. Needs to be removed.
  3. Reflections from the QPD and the PDA255 are not dumped. This causes noise. Bad.
  4. The QPD transimpedance should be reduced so that it can handle more light. I don't know what it has now, but its probably 10-100 kOhm.

We estimated that the power in the IMC is (1 W)*Finesse/pi = 500 W. The MC2 Transmission spec is < 10 ppm, so the power on the table is probably ~5 mW. Since the PDA255 has a transimpedance of 10 kOhm and a max output power of 10V, it can handle up to ~1 mW. Probably we can get the QPD to handle 4 mW.

Gautam, Steve  3-27

We measured MC2 transmitted power right at the uncoated window ~2.5 mW  The beam was just a little bigger than the meter.

Attachment 1: 20180326_201929.jpg
20180326_201929.jpg
  13704   Mon Mar 26 16:10:33 2018 KiraUpdatePEMfinal setup sketch

I made sketches of the final setup. There will be a box in the rack that contains both the heater circuit and the temperature sensor boards. One of them is in the loop while the other isn't. Instead of having many cables leading to the can, there will only be these three, though they can be made into a single wire. It will be connected to the can through a D-9 connector. The second attachment is what will be inside of the box, with all the major wires and components labeled.

-----

Edit: I've canged the layout to (hopefully) make the labels easier to read. I've also added in a cable to the ADC that reads out the voltage across the 1 ohm resistor. I also attached the circuit diagrams for the heater circuit and the temperature sensors. The one for the heater circuit was made by Kevin and I used the same design, except I have LM7818 and LM7918, since the 15V ones were not available at the time I made the circuit. 

In addition, all the wires leading to the can will all be part of one bundle of wires (I didn't clearly indicate it as such). There will be a total of 6 wires: two are needed for the wire to supply power to the heater and will have a LEMO connector on the rack end and two are needed for each temperature sensor, which will be attached to the board directly on the rack end. 

Also, we don't need two voltage regulators for each temperature circuit. We can just have one of each of LM7815 and LM7915 to supply +/- 15V to the boards.

Attachment 1: heater_1_new.png
heater_1_new.png
Attachment 2: heater_2_new.png
heater_2_new.png
Attachment 3: HeaterCircuit.pdf
HeaterCircuit.pdf
Attachment 4: temp_sensor.png
temp_sensor.png
  13703   Mon Mar 26 10:15:20 2018 ArijitUpdateIOOPMC and IMC re-locked

[Gautam, Arijit]

PMC and IMC re-aligned and re-locked. Both cavities are staying locked. Arm cavities are also locked.

  13702   Mon Mar 26 09:25:18 2018 SteveUpdateVACVM1 opened

CC1 old MKS cold cathode gauge randomly turns on- off. This makes software interlock close VM1 to protect RGA  So the closed off RGA region pressure goes up and the result is distorted RGA scan.

CC1 MKS gauge is disconnected and VM1 opened. This reminds me that we should connect our interlocks to CC1 Hornet Pressure gauge.

Quote:

Pumpdown 80 at 511 days and pd80b at 218 days

Valve configuration:  special vacuum normal, annuloses are not pumped at 3 Torr, IFO pressure 7.4e-6 Torr at vac envelope temp 22 +- 1C degrres

 

 

Attachment 1: CC4VM1.png
CC4VM1.png
  13701   Fri Mar 23 12:45:08 2018 KiraUpdatePEMtest setup

I fit the data that we got from the test. The time constant for the cooling came out to be about 4.5 hours. The error is quite large and we should add a low pass filter to the temperature sensor eventually in order to minimize the noise of the measurements.

Attachment 1: seis_fit.png
seis_fit.png
  13700   Fri Mar 23 12:00:20 2018 ranaUpdatePEMtest setup

we don't ever want to use our 16 kHz real time system for such low frequency action; its main purpose is for real-time controls, whereas we are OK with multiple seconds of delay in a thermal loop. The Python PID script is sufficient and highly reliable (after years of testing).

  13699   Thu Mar 22 17:47:16 2018 Angelina PanSummary Proposed QPD Optical Arrangement
Attachment 1: IMG_0869.jpg
IMG_0869.jpg
  13698   Wed Mar 21 21:13:44 2018 gautamUpdateIOOIMC noise budget

I've added two curves to the NB. Both are measured (with FET preamp) at the output of the demod board, with the LO driven at the nominal level by the Wenzel RF source pickoff (as it would be when the IMC is locked) and the RF input connected to the IMC REFL PD. For one curve, I simply closed the PSL shutter, while for the other, I left the PSL shutter open, but macroscopically misaligned MC2 so that there was no IMC cavity. So barring RFAM, there should be no PDH signal on the REFL PD, but I wanted to have light on there. I'm not sure if I understand the difference between these two curves though, need to think on it. Perhaps the IMC REFL PD's optical/electrical response needs to be characterized?

Quote:
 

Next curve to go on here is the demod board noise with the PSL shutter closed but the IMC REFL PD connected to the RF input (or maybe even better, have light on the PD, but macroscopically misalign MC2 so there is no 29.5MHz PDH signal), just to make sure there isn't anything funky going on there...

 

Attachment 1: IMC_RF_noise_calib.pdf
IMC_RF_noise_calib.pdf
  13697   Wed Mar 21 17:31:21 2018 gautamUpdateIOOMC error point calibration

I did a preliminary noise budget of the transmitted frequency noise of the IMC. Attachment #1 shows the NB. I'm going to use this opportunity to revisit my IMC modeling. Some notes:

  1. The blue, green and red traces are from my measurement of the voltage noise of the demod board with LO driven by Marconi, RF input terminated, measured using the FET preamp + SR785, and calibrated into Hz/rtHz using the number from the immediate preceeding entry.
  2. The grey trace is measured by closing the PSL shutter, manually engaging the POX11 whitening, and looking at the calibrated POX. This sensing noise is ~100x higher than the IMC curves - but I think this makes sense as for the IMC, I am measuring directly at the output of the demod board, whereas for POX, the signal goes through the whole whitening infrastructure first.
  3. The purple trace is the calibrated X-arm control signal converted to Hz/rtHz.
  4. I've only showed the region up to ~1kHz as this is where the excess noise was seen in the original ALS study that precipitated this whole investigation.

Conclusion: From this study, assuming my PDH discriminant calibration was correct, looks like IMC demod / POX11 demod electronics noises are not to blame (this surprises me since there were apparently so many things wrong on the demod board, and yet that wasn't the worst thing in the IMC chain it would seem frown). The POX11 photodiode "dark" noise is also not the problem I think, given the grey curve. Next curve to go on here is the demod board noise with the PSL shutter closed but the IMC REFL PD connected to the RF input (or maybe even better, have light on the PD, but macroscopically misalign MC2 so there is no 29.5MHz PDH signal), just to make sure there isn't anything funky going on there...

Quote:
 

Using this, I can now make up a noise budget of sorts for the IMC sensing.

 

Attachment 1: IMC_RF_noise_calib.pdf
IMC_RF_noise_calib.pdf
  13696   Wed Mar 21 15:52:45 2018 gautamUpdateIOOMC error point calibration

As discussed at the meeting, I decided to calibrate the MC error point into physical units of Hz/rtHz (a.k.a. the PDH discriminant). This is to facilitate the debugging of the hypothesized excess IMC sensing noise. I did this as follows.

  1. Trust the POX calibration that was last updated in Aug 2017.
  2. Hook up spare DAC channel (piped from LSC rack to 1X2) to IN2 of IMC CM board.
  3. Inject excitation into MC error point via "IN2" input of the common board. For an excitation of 30cts with the IN2 gain at -32dB, I was able to see a peak in the calibrated X arm control signal that was ~x10 above the nominal noise level around 150Hz without seeing any nonlinear coupling effects in the DTT spectrum (I'm assuming 150Hz is sufficiently above the UGF of the X arm locking loop such that no loop correction is necessary).
  4. Took a spectrum of the IMC error signal, teed off into the SR785 at the I output of the demod board with the same linewidth as the DTT spectrum.
  5. Confirmed that without any excitation,
  6. Did the math to make these two peaks line up. The resulting calibration is: 13kHz/Vrms.

Math details:

  • DTT peak height @ 150 Hz with Hanning window, 25 avgs = 1.97e-4 nm/rtHz (See Attachment #1).
  • X arm cavity length = 37.79m, using which the above number becomes 1.47 Hz/rtHz.
  • Peak height in SR785 spectrum with Hanning window, 25 avgs = 1.13e-4 Vrms/rtHz (See Attachment #2).
  • Dividing, we get 13kHz/Vrms.

Using this, I can now make up a noise budget of sorts for the IMC sensing.


gautam 20180327 4.30pm: I re-checked the PDH error signal calibration using the oscilloscope method. Attachment #3 shows the PDH I and Q error signals and also the output of the RF monitor port, during a TEM00 flash. This attachment should be compared to Attachment #2 of elog 12822, and the answer lines up quite well. From my Finesse model of the IMC, I calculated that the x-axis of the PDH horn-to-horn is ~12.3kHz. Comparing to the top row of Attachment #3, I get a PDH error signal calibration of ~12.4kHz/Vrms, which lines up well with the number quoted above. So I trust my calibration, and hence, the y-axis of my noise budgets in reply to this elog.

Attachment 1: IMC_PDHdisc_20180321.pdf
IMC_PDHdisc_20180321.pdf
Attachment 2: IMC_PDHdisc785_20180321.pdf
IMC_PDHdisc785_20180321.pdf
Attachment 3: IMC_oscope.pdf
IMC_oscope.pdf
  13695   Wed Mar 21 10:00:35 2018 steveUpdateGeneralprojector light bulb replaced

Light bulb replaced.

Quote:

Bulb went out ~10am today. Looks like the lifetime of this bulb was <100 days.

Steve: bulb is arriving next week

 

  13694   Tue Mar 20 22:44:45 2018 gautamUpdateIOOIMC loop checkup

After some persistence, I managed to get the mixers off.

  • Having gotten the mxiers off, I decided to temporarily solder on 50ohms between the LO pin pad and ground on the demod board and measure the RF signal levels in the LO chain with the active probe again.
  • Today, with this change, I confirmed that the ERA-5SM begins to saturate closer to the +19dBm advertised on its datasheet. So we need only 2dB of attenuation at the input to have 17dBm at the LO pin of the mixer, assuming 50ohm input impedance.
  • But this begs the question - what does minicircuits mean by a Level-YY mixer? Do they expect YY dBm delivered to a 50ohm load? Or do we need to supply YY dBm accounting for the dynamically changing input impedance of the mixer, as monitored by a high impedance probe?
  • I soldered on some new mixers (JMS-1H) I procured from Downs earlier today.
  • Re-installed the demod board in the eurocrate.

Unfortunately, the coherent noise between the arms persists so the sensing noise injection must be happening elsewhere. frownIMC seems to lock fine though so I'm leving the autolocker on

  13693   Tue Mar 20 21:08:03 2018 gautamUpdateIOOIMC loop checkup

This elog by koji inspired me to consider power supply as a possible issue.

The demod board receives +/-24V DC (which is regulated down to +/-15V DC by 7815/7915), and also +15V DC via the backplane. The ERA-5SM receives DC power from the latter (unregulated) +15V DC. I can't think of why this is the case except perhaps the regulators can't source the current the amp wants? In any case, it doesn't look feasible to change this by cutting any traces on the PCB to me. While I had the board out, I decided to replace the JMS-1H mixers in a last ditch effort to improve the demod board noise. Unfortunately I'm having trouble de-soldering these MCL components from the board. So for now, I'm leaving the demod board out, IMC unlocked. Work will continue tomorrow. 

  13692   Tue Mar 20 19:48:10 2018 gautamUpdatePEMtest setup

according to the temp sensor readout, which was ~-3.35V which corresponds to ~335K, the temperature of the can is now 60 deg C. This is a bit warm for my liking so i'm turning the heater current down to 0 now by writing 0 to C1:PEM-SEIS_EX_TEMP_CTRL

  13691   Tue Mar 20 16:56:01 2018 KiraUpdatePEMtest setup

The MOSFET was getting pretty hot, so I switched it out to a larger heat sink and it's fine now. I then used a function generator in place of the DAC to provide ~3.5V. I got the current in the circuit to 1.7A, which is as expected, since we have 24V input, the heater resistance is 12.5ohm and the resistor we are using is 1ohm, so 24V/(12.5+1)ohm = 1.7A. The temperature inside the can rose about 5 degrees in half an hour. The only issue now is the voltage regulators and OP amp inside the box get hot, though it doesn't seem to be dangerous. I switched the function generator input to a DAC and Gautam set it to 1.5V. If it works, then we'll leave this on overnight and work on the PID control tomorrow. I've attached images of the current heater circuit box when it is open and the new heat sink for the MOSFET.


gautam: we also tried incorporating the EPICS channels from the Acromag into the RTCDS so that we can implement PID control by using Foton. I tried doing this using the "EpicsIn" and "EpicsOut" blocks from CDS_PARTS. While the model recompiled smoothly, I saw no signals in the filter module i had connected in series with the EpicsIn block. So I just reverted c1pem to its original state and recompiled the model. Guess we will stick to python script PID reading EPICS channels to implement the PID servo.

Attachment 1: IMG_20180320_154516.jpg
IMG_20180320_154516.jpg
Attachment 2: IMG_20180320_145957.jpg
IMG_20180320_145957.jpg
  13690   Tue Mar 20 16:53:03 2018 gautamUpdateIOOIMC loop checkup

Re-measured the demod board noises after replacing the suspect ERA-5SMs, with LO driven by a marconi at the "nominal" level of 2.5dBm, and RF input terminated. Attachment #1 is the input referred voltage noise spectra. I used the FET low noise pre-amp box for this purpose. I cannot explain the shape of the spectra above 1kHz. I tried doing the measurement on a minicircuits mixer (non-surface mount) and found the shape to be flat throughout the SR785 span. Unclear what else could be going on in the demod board though, all the other components on it are passive (except the ERA-5SMs which were replaced). I considered adopting a PMC style demod setup where we do the demod using some separate Minicircuits Mixer+LowPass filter combo. But the RF flashes for the IMC monitored at the RFmon port are ~0.2Vpp, and so the RF input to the mixer is expected to be ~2Vpp. The minicircuits mixer selection guide recommends choosing a diode mixer with LO level at least 10dBm above the expected RF input signal level, and we don't have any standalone mixers that are >Level 7. I've asked Steve to package the aLIGO demod board in the meantime, but even that might not be a plug and play replacement as the IF preamp stage has ~120degrees phase lag at 1MHz, which is significantly higher than the existing board which just has a SCLF5 low pass filter after the mixer and hence has <45degrees phase lag at 1MHz.

Attachment 1: IMC_RF_noise.pdf
IMC_RF_noise.pdf
  13689   Mon Mar 19 23:44:00 2018 gautamUpdateIOOIMC loop checkup

[koji, gautam]

  1. I began my investigations by measuring the voltage noise of the demod board outputs with an SR560 (G=100) and SR785 in the audio band.
    • Measurement made with PSL shutter closed, LO input of demod board driven with the nominal level of ~2.5dBm, RF input terminated.
    • Motivation was to look for any noise features.
    • Expected noise level is ~2nV/rtHz (Johnson noise of 50ohm) since there are no preamp electronics post SCLF-5 LP filter on this board.
    • Attachment #1 shows the results of the measurement for a few scenarios. Spectra only shown for the I channel but the Q channel was similar. The LO=+5dBm curve corresponds to driving the input at 5dBm with a marconi, to see if the label of the nominal level being +5dBm had anything to it.
    • The arches above 1kHz seemed suspicious to me, so I decided to investigate further.
  2. Looking at the IMC Demod board schematic, I I saw that there were 2 ERA-5SMs in there which are responsible for amplifying the 29.5MHz signal which serves as the LO to the oscillators.
  3. I pulled the demod board out and tested it on the electronics workbench. Koji and I couldn't make sense of the numbers we were seeing (all measurements made with Agilent analyzer and active FET probe with 100:1 attentuator).
  4. We eventually concluded that the ERA-5SMs were not exhibiting the expected gain of ~20dB. So we decided to swap these out.
  5. This sort of measurement is not ironclad as the output of the ERA-5SM goes to the mixer whose input impedance is dynamically varying as the diodes are switching. So even after replacing the suspect amplifiers with new ones, we couldn't make the numbers jive.
  6. We suspected that the new amplifiers were getting saturated. The 3dB saturation point for the ERA-5SM is spec'ed as ~19dBm.
    • We "measured" this by varying the input signal level and looking for deviation from linearity.
    • We saw that there was ~1dB compression for ~13dBm output from the ERA-5SM (after correcting for all attenuators etc). But this number may not be accurate in the absolute sense because of the unknown input impedance of the mixer.
    • Moreover, looking at the spec sheet for the mixer, JMS-1H, we found that while it wasn't ideal to operate the mixer with the LO level a few dBm below the expected +17dBm, it probably wasn't a show stopper.
  7. So we figured that we need 10dB of attenuation between the "nominal" LO input level of 2.7dBm and the input of the demod board in order to keep the ERA-5SM in the linear range. This has now been implemented in the form of an SMA attenuator.
  8. IMC locked straight away. But I noticed that PC drive RMS level was unusually large.
  9. I found that by increasing the "IN1" gain of the CM board to 12dB (from 2dB) and the "VCO gain" to 10dB (from 7dB), I could recover a transfer function with UGF ~140kHz and PM ~30degrees (need more systematic and wider span measurements of this, and also probably need to optimize the crossover gains). See Attachment #2 for my quick measurement tonight.
  10. Updated mcup to reflect these new gains. Tested autolocker a few times, seems to work okay.
  11. While it presumably was a good thing to replace the faulty amplifiers and prevent them from saturating, this work has not solved the primary problem of excess frequency noise on the PSL.

It is not clear to me why installing an attenuator to prevent amplifier saturation has necessitated a 10dB increase in the IN1 gain and 3dB increase in the VCO gain. Initially, I was trying to compensate for the gain by increasing the FSS "Common Gain" but in that setting, I found an OLTF measurement impossible. The moment I enabled the excitation input to the CM board, the lock was blown, even with excitation amplitudes as small as -60dBm (from the Agilent network analyzer).

This may also be a good opportunity to test out one of the aLIGO style FET mixer demod boards (recall we have 2 spare from the 4 that were inside the ALS demod box). I'm going to ask Steve to package these into a 1U chassis so that I can try that setup out sometime. From a noise point of view, the aLIGO boards have the advantage of having a x100 preamp stage straight after the mixer+LPF. We may need to replace the lowpass filter though, I'm not sure if the one installed is 1.9MHz or 5MHz.

I've left an SR785 and AG4395 near 1X2 in anticipation of continuing this work tomorrow.

Unrelated to this work - seems like the WFS DC and RF offsets had not been set in a while so I reset these yesterday. The frequent model restarts in recent times may mean that we have to reset these to avoid using dated offset values.

Attachment 1: IMC_RF_noise.pdf
IMC_RF_noise.pdf
Attachment 2: IMC_OLTF_20180320.pdf
IMC_OLTF_20180320.pdf
  13688   Mon Mar 19 15:02:29 2018 gautamUpdateALSNoisy MC sensing

The working hypothesis, since the excess noise in single arm locks is coherent between both arms, the excess sensing noise is frequency noise in the IMC locking loop (sensing because it doesn't show up in MC_F). I've started investigating the IMC sensing chain, starting with the power levels of the RF modulation source. Recall that we had changed the way the 29.5MHz signal was sent to the EOM and demod electronics in 2017. With the handheld RF power meter, I measured 13.2dBm coming out of the RF distribution box (this is routed straight from the Wenzel oscillator). This is amplified to 26dBm by an RF amplifier (ZHL-2-S) and sent to the EOM, with a coupled 16dBm part sent to a splitter that supplies the LO signal to the demod board and also the WFS boards. Lydia made a summary of expected RF power levels here, and I too seem to have labelled the "nominal" LO level to the MC_REFL demod board as +5dBm. But I measured 2.7dBm with the RF power meter. But looking closely at the schematic of the splitting circuitry, I think for a (measured) 16.7dBm input to it, we should in fact expect around 3dBm of output signal. So I don't know why I labelled the "nominal" signal level as 5dBm.

Bottom line: we are driving a level 17 mixer with more like +14dBm (a number inferred from this marked up schematic) of LO, which while isn't great, is unlikely to explain the excess noise I think (the conversion loss just degrades by ~1dB). So I will proceed to check further downstream in the signal chain.

  13687   Mon Mar 19 14:39:09 2018 johannesConfigurationComputersc1auxex replacement

[gautam, johannes]

The temperature control output channel for the XEND seismometer wasn't working properly. The EPICS channel existed, could be written to and read from, but no physical voltage was observed on the (confirmed properly) wired connector.

The Acromag DAC that outputs this channel was completely spare in the original scheme and does not serve any other channels at the moment. We found it to be unresponsive to ping from the host machine (reminder: the Acromags are on their own subnet with IPs 192.168.114.xxx connected to the secondary ethernet adapter of c1auxex), while all others returned the ping just fine. The modules have daisy-chained ethernet connections, and the one Acromag unit behind the unresponsive one in the chain was still responding to ping and its channels were working, so it couldn't have been a problem with the (ethernet) cabling.

Gautam and I power-cycled the chassis and server, which resolved the issue. The channel is now outputting the requested voltage on the Out1 BNC connector of the chassis (front). When I was setting up the whole system and did frequent rebooting and IP-redefinitions I have seen network issues arise between server and Acromags. In particular, when changing the network settings server-side, the Acromags needed to reboot occasionally. So this whole problem was probably due to the recent server-swap, as the chassis had not been power-cycled since.

 

During the debugging we also found that the c1psl2 channels were not working. This was because I had overlooked to update the epics environment variables for the modbus path defined in /cvs/cds/caltech/target/c1psl2/npro_config.cmd from the local installation /opt/epics/ (which doesn't exist on the new server anymore) to the network location /cvs/cds/rtapps/epics-3.14.12.2_long/. This has been fixed and the slow diagnostic PSL channels are recording again.

  13686   Mon Mar 19 07:37:00 2018 Angelina PanSummary Proposed QPD Optical Arrangement

I am currently working on an optical arrangement consisting of a QPD that measures the fluctuations of an incoming HeNe laser beam that is reflected by a mirror. The goal is to add a second QPD to the optical arrangement to form a linear combination that effectively cancels out the (angular) fluctuations from the laser beam itself so that we can only focus on the fluctuations produced by the mirror.

In order to solve this problem, I have written a program for calculating the different contributions of the fluctuations of the HeNe laser and fluctuations from the mirror, for each QPD (program script attached). The goal of the program is to find the optimal combination of L0, L1, L2, and f2 that cancels the fluctuations from the laser beam (while retaining  solely the fluctuations from the mirror) when adding the fluctuations of QPD 1 and QPD 2 together. 

By running this program for different combinations of distances and focal lengths, I have found that the following values should work to cancel out the effects of the oscillations from the HeNe laser beam (assuming a focal length of 0.2 m for the lens in front of the original QPD):

  • L0 = 1.0000 m (distance from laser tube to mirror)
  • L1 = 0.8510 m (distance from mirror to lens in front of QPD 1)
  • L2 = 0.9319 m (distance from beamsplitter to lens in front of QPD 2)
  • f2 = 0.3011 m (focal length of lens in front of QPD 2)

Based on these calculations, I propose to try the following lens for QPD 2:

1’’ UV Fused Silica Plano-Convex Lens, AR-Coated: 350 - 700 nm (focal length 0.3011 m). https://www.thorlabs.com/newgrouppage9.cfm?objectgroup_id=6508

Attachment 1: angelinaCode.py.tar.bz2
  13685   Fri Mar 16 09:36:56 2018 SteveUpdateVACRGA scan at day 511,218d

Pumpdown 80 at 511 days and pd80b at 218 days

Valve configuration:  special vacuum normal, annuloses are not pumped at 3 Torr, IFO pressure 7.4e-6 Torr at vac envelope temp 22 +- 1C degrees

Quote:

pd80b rga scan at 175 day.  IFO pressure 7.3e-6 Torr-it

Condition: vacuum normal, annuloses not pumped. Rga turned on yesterday.

The rga was not on since last poweroutage Jan 2, 2018 It is warming up and outgassing Atm2

 

Attachment 1: RgaScan511d.png
RgaScan511d.png
Attachment 2: 08.png
08.png
  13684   Thu Mar 15 17:33:56 2018 KiraUpdatePEMtest setup

I have attached the setup I completed today. The metal box contains the heater circuit and the board for the temperature sensor is right above it. This is basically the same setup as before, but I've just packaged everything up neater. I expect to be able to perform the test tomorrow and begin implementing PID control. I still need a DAC input for the heater circuit and the temperature sensor is having some issues as well.

Attachment 1: IMG_20180315_172512.jpg
IMG_20180315_172512.jpg
  13683   Thu Mar 15 16:00:25 2018 Larry WallaceSummaryComputersCert renewal for NODUS

The cert for nodus has been renewed for another 2 years.

The following is the basic procedure for getting a new cert: (Note certs are only good for two years as of 2018)
openssl req -sha256 -nodes -newkey rsa:2048 -keyout nodus.ligo.caltech.edu.key -out nodus.ligo.caltech.edu.csr
Country Name (2 letter code) [AU]:US
State or Province Name (full name) [Some-State]:CaliforniaLocality Name (eg, city) []:Pasadena
Organization Name (eg, company) [Internet Widgits Pty Ltd]:California Institute of Technology
Organizational Unit Name (eg, section) []:LIGO
Common Name (eg, YOUR name) []:nodus.ligo.caltech.edu

Leave the e-mail address, challenge password and optional company name blank. A new private key will be generated.
chown root nodus.ligo.caltech.edu.key
chgrp root nodus.ligo.caltech.edu.key
chmod 0600 nodus.ligo.caltech.edu.key

The nodus.ligo.caltech.edu.csr file is what is sent in for the cert.
This file should be sent to either ryan@ligo.caltech.edu or security@caltech.edu and copy wallace_l@ligo.caltech.edu.

A URL llink with the new cert to be downloaded will be sent to the requestor.

Once the files are downloaded, the new cert and intermediate cert, they can be copied and renamed.

The PEM-encoded host certificate by itself is saved at:

  /etc/httpd/ssl/nodus.ligo.caltech.edu.crt

The nodus.ligo.caltech.edu.key file should be in the same directory or whichever directory is indicated in the ssl.conf located in /etc/httpd/conf.d/  directory.

httpd will need to be restarted in order for it to see the new cert.

 

  13682   Wed Mar 14 23:58:30 2018 johannesConfigurationComputersc1auxex replacement

I replaced the borrowed server with the permanent one today. Before Removing the current server, Before, I performed several additional preparations:

  • Updated Chiara hostables to IP 192.168.113.48 for c1auxex
  • apt-get install procserv
  • copied ETMXaux2.* files in /cvs/cds/caltech/target/c1auxex2 to ETMXaux.* and changed references from /opt/rtcds/epics (which was a local directory on c1auxex2) to /cvs/cds/rtapps/epics-3.14.12.2_long in the copied files
  • Added instruction
    Environment="LD_LIBRARY_PATH=/cvs/cds/rtapps/epics-3.14.12.2_long/base/lib/linux-x86_64:/cvs/cds/rtapps/epics-3.14.12.2_long/modules/modbus/lib/linux-x86_64/:/cvs/cds/rtapps/epics-3.14.12.2_long/modules/asyn/lib/linux-x86_64"
    to /etc/systemd/system/modbusIOC.service  (required for burtwb dependencies)

Then I replaced the server:

  1. IFO was in LSC mode with both arms locked
  2. Backed up ETMX alignment using save feature in IFOalign screen
  3. Disengaged LSC mode
  4. Shut down ETMX watchdog
  5. Disconnected ETMX satellite box
  6. Shut down c1auxex2 and c1auxex
  7. Performed the server swap
  8. Booted c1auxex
  9. Made sure EPICS channels were back online and channel defaults were restored
  10. Reconnected satellite box
  11. Turned on watchdog
  12. Turned on OpLevs
  13. Engaged LSC mode -> both arms were instantly locked

I returned c1auxex2 to Larry, who needed it back asap because of some hardware failure

Steve: Acromag XT1221 ordered 3-15-18

  13681   Tue Mar 13 20:03:16 2018 johannesConfigurationComputersc1auxex replacement

I assembled the rack-mount server that will long-term replace c1auxex, so we can return the borrowed unit to Larry.

SUPERMICRO SYS-5017A-EP Specs:

  • Intel Atom N2800 (2 cores, 1.8GHz, 1MB, 64-bit)
  • 4GB (2x2GB) DDR3 RAM
  • 128 GB SSD

IMG_20180313_105154890.jpg      IMG_20180313_133031002.jpg

I installed a standard Debian Jessie distribution, with option LXDE for minimal resource usage. Steps taken after fresh install

  1. Give controls sudo permission: usermod -aG sudo controls
  2. mkdir /cvs/cds
  3. apt-get install nfs-common
  4. Added line "chiara:/home/cds              /cvs/cds        nfs     rw,bg,nfsvers=3" to end of /etc/fstab
  5. Configured network adapter in /etc/network/interfaces
            iface eth0 inet static
            address 192.168.113.48
            netmask 255.255.255.0
            gateway 192.168.113.2
            dns-nameservers 192.168.113.104 131.215.125.1 131.215.139.100
            dns-search martian

    I first assigned the IP 192.168.113.59 of the original c1auxex, but for some reason my ssh connections kept failing mid-session. After I switched to a different IP the disruption no longer happened.
  6. Add lines "search martian" and "nameserver 192.168.113.104" to /etc/resolv.conf
  7. apt-get install openssh-server
    At this point the unit was ready for remote connections on the martian network, and I moved it to the XEND.
  8. Added lines to /home/controls/.bashrc to set paths and environment variables:
    export PATH=/cvs/cds/rtapps/epics-3.14.12.2_long/base/bin/linux-x86_64:/cvs/cds/rtapps/epics-3.14.12.2_long/extensions/bin/linux-x86_64:$PATH
    export HOST_ARCH=linux-x86_64
    export EPICS_HOST_ARCH=linux-x86_64
    export RPN_DEFNS=~/.defns.rpn
    export LD_LIBRARY_PATH=$LD_LIBRARY_PATH:/cvs/cds/rtapps/epics-3.14.12.2_long/base/lib/linux-x86_64:/cvs/cds/rtapps/epics-3.14.12.2_long/modules/modbus/lib/linux-x86_64/:/cvs/cds/rtapps/epics-3.14.12.2_long/modules/asyn/lib/linux-x86_64
  9. apt-get install libmotif-common libmotif4 libxp6 (required to run burtwb utility)

The server is ready to take over for c1auxex2 and does not need any local epics compiled, since it can run the 3.14.12.2_long binaries in /cvs/cds.

Attachment 1: IMG_20180313_105154890.jpg
IMG_20180313_105154890.jpg
Attachment 2: IMG_20180313_133031002.jpg
IMG_20180313_133031002.jpg
  13680   Mon Mar 12 23:57:31 2018 gautamUpdateALSNoisy POX

[kevin, gautam]

Kevin suggested I shouldn't be so lazy and test the POY spectrum as well. So we moved the timing card back to c1iscey, went through the usual dance of vertex machine reboots, and then got both single arm locks going. Attached spectrum shows that both POX and POY are noisy. I'm not sure what has changed that could cause this effect. The fact that both POX and POY appear uniformly bad, but that there is no coherence with MC_F, suggests to me that perhaps this has something to do with the work I did with Koji w.r.t. the power situation at the LSC rack. But we just checked that

  1. All the demod board front panel LED indicators are green.
  2. Marconi and all RF amplifier boxes are on (but we didn't actually measure any RF power levels yet tonight).
  3. We checked the KEPCO power supplies in the little cabinet along the Yarm, and all of them are reporting the correct voltages/currents as per Steve's (recently updated) labels.
  4. Checked the expansion chassis at the LSC rack for any red lights, there were none.

Another observation we made: note the huge bump around 70Hz in both arm control signals. We don't know what the cause of this is. But we occassionally noticed harmonics of this (i.e. 140, 210 Hz etc) appear in the control signal spectra, and they would grow with time - eventually, the X arm would lose lock (though the Y arm stayed locked).

I'm short on ideas for now so we will continue debugging tomorrow.


Unrelated to this work: Kevin reminded me that the high-pitched whine from the CRT TVs in the control room (which is apparently due to the flyback transformer) is DEAFENING. It's curious that the "chirp" to the eventual 15kHz whine is in opposite directions for the QUAD CRTs and the single display ones. Should be a Ph6 experiment maybe.


Update 2:30pm Mar 13: The furthest back I seem to be able to go in time with Frames is ~Jan 20 2018. Looking for a time when the arms were locked from back then, it seems like whatever is responsible for a noisy POX and POY was already a problem back in January. See Attachment #2. So it appears that the recent work at 1Y2 is not to blame...

Attachment 1: NoisyPOXandPOY_20180312.pdf
NoisyPOXandPOY_20180312.pdf
Attachment 2: noisyPOX_Jan2018.pdf
noisyPOX_Jan2018.pdf
  13679   Mon Mar 12 22:08:31 2018 gautamUpdateALSNoisy POX

[kevin, gautam]

we tested my noisy POX hypothesis tonight. By locking the single arm with POX, the arm length is forced to follow PSL frequency, which is itself slaved to IMC length. From Attachment #1, there is no coherence between the arm control signal and MC_F. This suggests to me that the excess noise I am seeing in the arm control signal above 30 Hz is not originating from the PSL. It also seems unlikely that at >30Hz, anything mechanical is to blame. So I am sticking with the hypothesis that something is wonky with POX. For reference, a known "normal" arm control signal spectrum looks like the red curve in this elog.

 

Attachment 1: NoisyPOX_20180312.pdf
NoisyPOX_20180312.pdf
  13678   Mon Mar 12 13:58:37 2018 gautamUpdateGeneralprojector light bulb blown

Bulb went out ~10am today. Looks like the lifetime of this bulb was <100 days.

Steve: bulb is arriving next week

Quote:

Bulb  is replaced.

  13677   Fri Mar 9 20:35:41 2018 Udit KhandelwalSummaryGeneralSummary 2018/03/09

1. Optical Table Layout 

I had discussed with Koji a way to record coordinates of optical table equipments in a text file, and load to solidworks. The goal is to make it easier to move things around on the table in the CAD. While I have succeeded in importing coordinates through txt files, there is still a lot of tediousness in converting these points into sketches. Furthermore, the task has to be redone everytime a coordinate is added to or changed in the txt file. Koji and I think that this can all be automated through solidworks macros, so I will explore that option for the next two weeks.

2. Vacuum Chamber CADs 

Steve will help find manufacturing drawings of the BS chamber. I have completed the ETM chambers, while the ITM ones are identical to them so I will reuse parts for the CAD. 

  13676   Fri Mar 9 12:59:53 2018 KiraUpdatePEMADC noise measurement

[Kira, Gautam]

I ceated a simple circuit that takes in 15V and outputs precisely 5V by using a 12V voltage regulator LM7812 and an AD586 that takes the output of the voltage regulator and outputs 5V (attachment 1). We plugged this into the slow channel and will leave it running for a few hours to see if we still have the fluctuations we observed earlier and also fit the noise curve. We'll also test the fast channel later as well. Attachment 2 shows the setup we have in the lab, with the red and white cable plugged into the +15V power supply and the red and black cable connected to the slow channel.

Attachment 1: IMG_20180309_114345.jpg
IMG_20180309_114345.jpg
Attachment 2: IMG_20180309_125153.jpg
IMG_20180309_125153.jpg
  13675   Fri Mar 9 01:07:01 2018 gautamUpdateALSFirst look at new ALS electronics

[koji, gautam]

I was going to head out but then it occurred to me that I could do another simple test, which is to try and lock the X arm on ALS error signal (i.e. actuate on MC length to keep the beat between EX laser and PSL fixed, while the EX frequency is following the Xarm length). Comparing the in loop (i.e. ALS) error signal with the out-of-loop sensor (i.e. POX), it seems like POX is noisy. The curves were lined up by eye, by scaling the blue curve to match the red at the ~16Hz peaks. This supports my hypothesis in the previous elog. On the downside, could be anything. Electronics in the POX chain? The demod unit itself? Will look into it more tomorrow..

As an aside, controlling the arm with ALS error signal worked quite well, and the lock was maintained for ~1 hour.

Attachment 1: ALS_vs_POXnoise.pdf
ALS_vs_POXnoise.pdf
  13674   Thu Mar 8 23:50:27 2018 gautamUpdateALSFirst look at new ALS electronics
  • Locked single arms, dither aligned, and saved offsets to EPICS (slow) sliders in anticipation of having to reboot all vertex FEs.
  • Shutdown ETMY watchdog, stopped all models on c1iscey, and shutdown that frontend.
  • Walked down to Y-end, powered of c1iscey expansion chassis, and removed the ADC adaptor card.
  • Stopped all models on c1lsc. Shutdown watchdog on all optics in anticipation of c1sus model failing. Shutdown the c1lsc frontend.
  • Powered off the c1lsc expansion chassis. Installed the borrowed adapter card from c1iscey in c1lsc expansion chassis. Connected it to the "spare" ADC card Koji and I had installed in c1lsc expansion chassis last Wednesday.
  • Connected differential output of demod board to differential input of AA chassis. Connected SCSI connector from output of AA chassis to the newly installed adapter card.
  • Powered the c1lsc expansion chassis back on. Then powered c1lsc FE on.
  • Walking back out to the control room, I saw that all vertex FEs had crashed. I had to go back in and hard-reboot c1sus.
  • Before bringing back any models, I backed up the existing c1lsc model, and then modified c1x04 and c1lsc to use the newly acquired ALS signals for the X arm ALS signal chain.
  • Restarted all vertex FE models. Everything came back up smooth. DC light is still red on c1oaf but I didn't bother trying to rectify it tonight for these tests.
  • Reset appropriate LSC offsets with PSL shutter closed. Locked X arm on IR. Reset phase tracker servo gain for X arm ALS. Engaged slow temperature servo on EX laser.

Then I looked at  the spectrum, see Attachment #1. Disappointingly, it looks like the arm PDH servo is dominating the noise, and NOT unsuppressed EX laser frequency noise,. Not sure why this is so, and I'm feeling too tired to debug this tonight. But encouragingly, the performance of the new ALS signal chain looks very promising. Once I tune up the X arm loop, I'm confident that the ALS noise will be at least as good as the reference trace.

I am leaving c1iscey shutdown until this is fixed. So ETMY is not available for the moment.

Random factoid: Trying to print a DTT trace with LaTeX in the label text on pianosa causes the DTT window to completely crash - so if you dont save the .xml file, you lose your measurement.

Quote:

I made a LISO fit of the measured TF of the daughter board, so that I can digitally invert the daughter board whitening. Results attached. (Inverse) Filters have been uploaded to the ALS X Foton filter banks.

 

Attachment 1: BeatMouth_OOL.pdf
BeatMouth_OOL.pdf
  13673   Thu Mar 8 19:38:37 2018 gautamUpdateALSdigital unwhitening of daughter board

I made a LISO fit of the measured TF of the daughter board, so that I can digitally invert the daughter board whitening. Results attached. (Inverse) Filters have been uploaded to the ALS X Foton filter banks.

Attachment 1: TFfit.pdf
TFfit.pdf
  13672   Thu Mar 8 18:15:42 2018 gautamUpdateGeneralCDS recovery after work at LSC rack

I was forced into a simultaneous power-cycle rebooting of the three vertex FEs just now. I took the opportunity to completely disconnect the c1sus expansion chassis from all power and then restart it.

Everything is back up right now, and the weird timing issues I noticed in the sus model seem to be gone now (I'll need a longer baseline to be sure and I'll post a trend of the CPU timing tomorrow). It's disconcerting that apparently the only way to get everything back up and running is the nuclear option of power-cycling all FE related electronics. I was considering borrowing an ADC adapter card from the Y end and measuring the calibrated IR ALS noise with the digital system, but if I'm going to have to go through this whole dance each time I do a model recompile on c1lsc (which I'm going to have to in order to get the extra ADC recognized), I'm wondering if it's just better to wait till we get the new adapter cards we ordered. I think I'm going to work on tuning the input coupling into the fiber at EX in the next couple of days instead.

Quote:
 

Seems like some high level voodoo indecision.


Edit 330pm: The model just crashed again. dmesg rather unhelpfully just says "ADC timeout". Unclear how to debug further. See Attachment #3.

 

  13671   Thu Mar 8 15:23:16 2018 gautamUpdateElectronicsNew DC power ports at c1lsc

[Koji, Gautam]

Yesterday, we installed some new DIN rail connectors at the LSC rack to provide 3 new outputs each for +24V DC and -24V DC. The main motivation was to facilitate the installation and powering of the differential receiving AA board. The regulators used inside the 1U chassis actually claims a dropout voltage of 0.5V and outputs 14V nominally, so a +/-15V DC supply would've perhaps been sufficient, but we decided to leave a bit more margin, and unfortunately, there are no +/-18V DC KEPCO linear power supplies to the LSC rack. Procedure:

  1. Prepared a bunch of DIN rail connectors with tinned, daisy-chained wires in the office area. Checked continuity and isolation with DMM.
  2. Checked that the two Sorensens at the bottom of the LSC rack were powering the RF distribution box and nothing else at the LSC rack.
  3. Walked over to the little rack housing all the KEPCO DC power supplies that supply DC voltages to the LSC rack. After checking that the labelled voltage and current values were correct, we turned them off, first +/-5V, then +/-15V (2 sets), and finally +/-24V.
  4. Installed the pre-assembled DIN connectors on the side rail at the LSC rack (we had to remove the side panel for the rack to do this work).
    • We used the ports supplying power to the ALS 1U demod chassis (+/-24V DC) to tap these voltages to our newly installed connectors.
    • The interconnecting wires are rather thick gauge, and especially for the ground wire, we found it impossible to push in our tap-off wire into the "correct/hot" side of the DIN blocks. So we had to use the other side instead. I'll upload a picture shortly which will make this more clear.
    • Checked continuity and isolation with DMM.
    • Turned the KEPCOs back on in reverse order to how they were turned off.
    • Measured voltages on the hot side of the DIN blocks, confirmed that they were as expected.
  5. Prepared a 12AWG aLIGO style power cable to connect to the 1U chassis. A reel of this cabling, with yellow shielding, is located ~halfway along under the EW arm. Koji prepared the actual connector and housed it in a DSUB shell as per aLIGO wiring color scheme.
  6. Installed the power cabling to one set of our 3 newly installed +/-24V DC power supplies.
  7. Inserted fuses into the hot DIN blocks, measured voltage at connector end of our newly installed power cable. At first, I forgot to check if the fuse blocks had fuses inside, but after this was rectified, voltages were as expected yes.

The c1lsc frontend models crashed for some reason during this procedure. Now the c1sus frontend model is also behaving weirdly. It is unclear to me if/how this work would have led to these problems, but the temporal correlation (but not causation?) is undeniable.

ELOG V3.1.3-